# Agile Decision-Making and Safety-Critical Motion Planning for Emergency Autonomous Vehicles

### v3

@@ III-D LSGM Framework
1:Initialization
2:Formulate $\text{Formulate nodes}\ \mathcal{N},$ Construct graph $\text{Construct graph connections}\ \mathcal{G}_{c}$
3:$n_{\text{start}}\leftarrow$ CurrentNode(), depth $d=1$
4:current traversal path $p_{c}$ is empty
5:visited level record $\mathcal{V}$ is empty
6:Gap Maginitude Judgement
7:$\mathcal{N}_{e}\leftarrow\text{GapMagnitudeJudge($\mathcal{N}$)}$
8:$\mathcal{N}_{e^{-}},\mathcal{G}_{ce^{-}}\leftarrow\text{UpdateGraphAndNodes}(\mathcal{N},\mathcal{N}_{e},\mathcal{G}_{c})$
9:LSGM Decision-Making
10:while True do
11:     $n_{\text{long}}\leftarrow\text{LongTermEfficiency}(\mathcal{N}_{e^{-}})$
12:     if $n_{\text{long}}==n_{\text{start}}$ then
13:         $n_{\text{target}}$ = $n_{\text{start}}$
14:         return $n_{\text{target}}$
15:     $\mathcal{P}_{\text{all}}\leftarrow\text{C-DFS}(\mathcal{N}_{e^{-}},\mathcal{G}_{ce^{-}},p_{c},\mathcal{V},n_{\text{long}},n_{\text{start}},d)$
16:     if $\mathcal{P}_{\text{all}}$ is not empty then
17:         $\mathcal{P}_{s}\leftarrow\text{FindShortestPath}(\mathcal{P}_{\text{all}})$
18:         if $|\mathcal{P}_{s}|==1$ then
19:              $n_{\text{target}}=\mathcal{P}_{s}.\text{first}$
20:         else
@@
22:         return $n_{\text{target}}$
23:     else
24:         $\mathcal{N}_{e^{-}},\mathcal{G}_{ce^{-}}\leftarrow\text{UpdateGraphAndNodes}(\mathcal{N}_{e^{-}},\mathcal{G}_{ce^{-}})$
IV Motion Planning Framework
@@
IV-A Constraints States
Constraint states represent different driving behaviors by defining the specific constraints required for each behavior. In this subsection, three types of constraint state are introduced: the lane-changing state, the lane-keeping state, and the lane-probing state. These states will be utilized to formulate constraints for the optimization problem.
Lane-keeping State. When the desired vehicle group is the current group, the ego vehicle maintains the current lane. For lane keeping, the required constraints are of two types: longitudinal constraints, for constraining the leading and following vehicles in the current lane longitudinally; and lateral constraints, for constraining vehicles around the ego vehicle within a certain range on adjacent lanes laterally.
Lane-probing State. If the desired vehicle group differs from the current vehicle group, it indicates support for a lane change. When the ego vehicle aims to move to a specific vehicle group, the primary interaction is with the follower in the desired group. The lane-probing state emerges when the ego vehicle has not yet attained a spatial advantage over the follower in the desired group, implying the need for the ego vehicle to proactively explore opportunities for merging into the desired group. The ego vehicle will continue to actively probe forward until it achieves the spatial condition $s_{ego}-s_{f}^{d}\geq\frac{l}{2}$, indicating that a sufficient spatial advantage has been gained. For lane-probing, the required constraints include longitudinal constraints for the following vehicle and ellipse constraints for the leading vehicle in the current group, in addition to lateral constraints. These ellipse constraints, described as linearized DHOCBF constraints in Section 3, enable the vehicle to safely probe forward for a spatial advantage while facilitating a smooth transition between lane-probing and lane-changing.
@@ IV-A Constraints States
IV-B Constraints Formulation
As mentioned in Section IV-A, different states are associated with specific constraints. This subsection elaborates on the different constraints formulation, including longitudinal constraints, lateral constraints, ellipse constraints, safety boundary constraints, and actuator constraints.
Longitudinal Constraints. To ensure the safety of the ego vehicle longitudinally, it is crucial to maintain appropriate distance from both the leading and following vehicles. Also, during lane-changing, longitudinal constraints on the desired group’s leader help ensure smooth adaptation. To achieve this, we adopt a linear DCBF aimed at defining a safety distance barrier. The function is mathematically formulated as
$h_{lon}(\mathbf{x})=|s-s_{i}|-T_{h}v_{x}-d_{0},$ (21)
where $s_{i}$ indicates the spacing of the surrounding vehicle with $i\in\{cl,cf,dl\}$. To be specific, $cl$ stands for the leader of the current group, $cf$ represents a follower in the current group, and $dl$ is the leader of the desired group. $T_{h}$ and $d_{0}$ denote the time headway and the least distance respectively. The longitudinal constraints can hence be constructed by (15) as follows
$\Delta h_{lon}^{i}(\mathbf{x}_{k},\mathbf{u}_{k})\geq-\gamma_{lon}^{i}h_{lon}^{i}(\mathbf{x}_{k})-\epsilon_{lon}^{i},$ (22)
@@ IV-B Constraints Formulation
$h_{lat}^{j}(\mathbf{x})=|e_{y}-e_{y}^{proj}|-w-d_{lat},\quad j\in\mathcal{J},$ (23)
where $\mathcal{J}=\{j\mid e_{y}^{\text{proj}}\in\text{ROI}\}$ , $e_{y}^{proj}$ denotes the deviation distance projected on the track centerline of the ego vehicle. Moreover, $w$ is the width of a vehicle and $d_{lat}$ is a positive constant. The lateral constraints can then be formed the same way as (15):
$\Delta h_{lat}^{j}(\mathbf{x}_{k},\mathbf{u}_{k})\geq-\gamma_{lat}^{j}h_{lat}^{j}(\mathbf{x}_{k})-\epsilon_{lat}^{j},$ (24)
@@
Ellipse Constraint. During lane-changing, we adopt linearized DHOCBF to formulate the constraint of leader in the current vehicle group or follower in the desired vehicle group. Different from [33], which used it for static obstacle avoidance and regarded obstacle as a circle, we apply it to dynamic traffic scenarios and adopt elliptical obstacles in the Frenet coordinate system.
@@
where the $s_{o}$ and $e_{yo}$ represent the traveled distance and lateral error of the considered ellipse border respectively.
@@
$L(s,e_{y},\lambda)=(s-s_{o})^{2}+(e_{y}-e_{y_{o}})^{2}$
$+\lambda\left(\frac{(s-s_{o})^{2}}{a^{2}}+\frac{(e_{y}-e_{y_{o}})^{2}}{b^{2}}\right),$ (26)
@@
The linearized DHOCBF is the tangent of the ellipse based on the closet point on the ellipse, the detailed formulation is as follows
@@
where for the follower in the desired vehicle group, the center of the ellipse is at the center of the front edge, while for the leader in the current vehicle group, the center of the ellipse is at the center of the rear edge.
Figure 6: Constraint mechanisms under different constraint states. In the LK state, the optimization problem includes longitudinal DCBF constraints for the current lane and lateral DCBF constraints for adjacent lanes. In the LP state, the leading vehicle in the current lane is transformed into a linearized DHOCBF constraint to better accommodate potential lane change maneuvers while actively exploring the feasible space. In the LC state, the rear vehicle in the desired group and the front vehicle in the current group are the main interaction targets. Linearized DHOCBF ensures lane change safety and expands the feasible region for lane changes.
@@
$-\lambda_{b}-\epsilon_{b}\leq e_{y}\leq\lambda_{b}+\epsilon_{b},$ (28)
where $\lambda_{b}$ is a small positive constant and $\epsilon_{b}$ is the slack variable for safety boundary constraints. Safety boundary constraints restrict the lateral error in a certain range centered on the centerline, if the ego vehicle goes far from the centerline, its cost increases due to a penalty on $\epsilon_{b}$. In short, we describe this constraint as follows:
$H_{S}(\mathbf{x})\leq\epsilon_{b}.$ (29)
@@
IV-C Motion Planner Formulations
We employed MPC to establish a finite-time convex optimization problem. Based on three different constraint states, we offer three optimization formulations. For simplicity, we abbreviate the three constraint states as LK (Lane-keeping), LP (Lane-probing), and LC (Lane-changing) optimizations. Fig. 6 provides a concise visual representation of the constraints associated with each of these states, illustrating how the optimization problem adapts to different driving conditions.
@@ IV-C Motion Planner Formulations
$\underset{\mathbf{u}_{t}:t+N-1|t,d_{1},d_{2}:t+N-1}{\arg\min}\sum_{k=0}^{N-1}||\mathbf{u}_{t+k}||_{Q}+||\dot{\mathbf{u}}_{t+k}||_{P}$
@@
$\underset{\mathbf{u}_{t}:t+N-1|t,d_{1},d_{2}:t+N-1}{\arg\min}\sum_{k=0}^{N-1}||\mathbf{u}_{t+k}||_{Q}+||\dot{\mathbf{u}}_{t+k}||_{P}$
@@
$\underset{\mathbf{u}_{t}:t+N-1|t,d_{1},d_{2}:t+N-1}{\arg\min}\sum_{k=0}^{N-1}||\mathbf{u}_{t+k}||_{Q}+||\dot{\mathbf{u}}_{t+k}||_{P}$
@@
Among these three optimization formulations, (31b), (32b), (33b) denote the constraints imposed by system dynamics. The vector $\mathbf{x}_{ref}$ represents the desired values for $v_{x}$, $w$, $v_{y}$, along with the desired values for $e_{y}$ and $e_{\psi}$, which are set to 00. This setting arises because, in Frenet coordinates, all points are projected onto the centerline that the ego vehicle is intended to follow, thereby making the desired deviation distance and heading angle error equal to zero. $p(\mathbf{x}_{t+N_{p}|t})$ is the terminal cost, where $S$ signifies the weight attributed to the terminal cost. Drawing upon the method described by [32] that employs a control Lyapunov function (CLF) as the terminal cost, we utilize the squared error $e_{\psi}^{2}$ as the terminal cost. $||\mathbf{u}_{t+k}||$, $||\dot{\mathbf{u}}_{t+k}||$, and $||\mathbf{x}_{t+k}-\mathbf{x}_{ref}||$ serve as stage costs, with $Q$, $P$, and $R$ representing the weights for the control inputs, input derivatives, and quadratic terms of tracking $\mathbf{x}_{\text{ref}}$, respectively. In particular, the weight matrix $R$ for $v_{x}$, $e_{\psi}$, and $e_{y}$ is designed to increase over time. This time-increasing strategy helps to prevent the generation of excessively large steering angles at the beginning of the trajectory.
V Experiment
In this section, we perform comprehensive simulations to evaluate the proposed system IDEAM. In section V-A, we introduce the simulation background about scenarios and parameters. In section V-B, we present the simulation benchmark and metrics. Section V-C gives the simulation results and section V-D shows a detailed case study of simulations.
V-A Scenario and Parameters Description
@@
Simulation Scenarios. The simulation scenario occurs on a continuous closed-loop path that includes both curves and straight sections, challenging the autonomous system to maintain consistent performance across the varying road geometry. The surrounding vehicles are randomly generated within the scene, with varying initial positions and desired velocities, ensuring a diverse set of interactions. The longitudinal motion of surrounding vehicles is simulated using the intelligent driver model (IDM) [37], while lateral control is achieved through a proportional-integral-derivative (PID) controller, the entire scenario takes place within a multi-lane environment, where the curvature of the bends varies along the path, reaching a maximum of approximately $0.1$. The dynamics of the scenario are updated at each step with a discretization interval of $\Delta t=100$ $ms$, following a kinematic model that dictates vehicle movements as outlined below,
$\dot{x}$ $=v\cos(\psi+\beta),$ (34a)
$\dot{y}$ $=v\sin(\psi+\beta),$ (34b)
$\dot{\psi}$ $=\frac{v}{l_{r}}\sin\beta,$ (34c)
$\dot{v}$ $=a,$ (34d)
$\beta$ $=\tan^{-1}\left(\frac{l_{r}}{l_{f}+l_{r}}\tan\delta_{f}\right).$ (34e)

### v4

@@ IV-D LSGM Framework
  • Complexity Analysis: The LSGM algorithm extends C-DFS by introducing an iterative decision-making process that evaluates long-term efficiency, executes C-DFS, and refines the decision space through graph updates. The main computational cost arises from the while loop, which iteratively performs efficiency evaluation ($O(N)$), C-DFS ($O(\min(P,b^{d},N+E))$), and graph updates ($O(N+E)$) until a valid decision is found. In the worst case, with up to $O(N)$ iterations and ineffective pruning, the overall complexity is $O(N(N+E))$. However, empirical results show that pruning significantly reduces the search space, resulting in an effective complexity closer to $O(NP)$.
@@
1:Input: nodes  $\mathcal{N}$, constructed graph $\text{constructed graph connections}\ \mathcal{G}_{c}$
2:         current node $n_{\text{start}}$, depth $d=1$
3:         empty traversal path $p_{c}$, empty level record $\mathcal{V}$
4:Output: target path $n_{\text{start}}$
5:Gap Maginitude Judgement
6:$\mathcal{N}_{e}\leftarrow\text{GapMagnitudeJudge($\mathcal{N}$)}$
7:$\mathcal{N}_{e^{-}},\mathcal{G}_{ce^{-}}\leftarrow\text{UpdateGraphAndNodes}(\mathcal{N},\mathcal{N}_{e},\mathcal{G}_{c})$
8:LSGM Decision-Making
9:while True do
10:     $n_{\text{long}}\leftarrow\text{LongTermEfficiency}(\mathcal{N}_{e^{-}})$
11:     if $n_{\text{long}}==n_{\text{start}}$ then
12:         $n_{\text{target}}$ = $n_{\text{start}}$
13:         return $n_{\text{target}}$
14:     $\mathcal{P}_{\text{all}}\leftarrow\text{C-DFS}(\mathcal{N}_{e^{-}},\mathcal{G}_{ce^{-}},p_{c},\mathcal{V},n_{\text{long}},n_{\text{start}},d)$
15:     if $\mathcal{P}_{\text{all}}$ is not empty then
16:         $\mathcal{P}_{s}\leftarrow\text{FindShortestPath}(\mathcal{P}_{\text{all}})$
17:         if $|\mathcal{P}_{s}|==1$ then
18:              $n_{\text{target}}=\mathcal{P}_{s}.\text{first}$
19:         else
@@
21:         return $n_{\text{target}}$
22:     else
23:         $\mathcal{N}_{e^{-}},\mathcal{G}_{ce^{-}}\leftarrow\text{UpdateGraphAndNodes}(\mathcal{N}_{e^{-}},\mathcal{G}_{ce^{-}})$
V Motion Planning Framework
@@
V-A Constraints States
Constraint states represent different driving behaviors by defining the specific constraints required for each behavior. In this subsection, three types of constraint states are introduced: the lane-changing state, the lane-keeping state, and the lane-probing state. These states will be utilized to formulate constraints for the optimization problem.
Lane-keeping State. When the desired vehicle group is the current group, the ego vehicle remains in the current lane. For lane keeping, the required constraints are of two types: longitudinal constraints, for constraining the leading and following vehicles in the current lane longitudinally; and lateral constraints, for constraining vehicles around the ego vehicle within a certain range on adjacent lanes laterally.
Lane-probing State. If the desired vehicle group differs from the current vehicle group, it indicates support for a lane change. When the ego vehicle aims to move to a specific vehicle group, the primary interaction is with the follower in the desired group. The lane-probing state emerges when the ego vehicle has not yet attained a spatial advantage over the follower in the desired group, implying the need for the ego vehicle to proactively explore opportunities for merging into the desired group. The ego vehicle will continue to actively probe forward until it achieves the spatial condition $s_{e}-s_{f}^{d}\geq\frac{l}{2}$, indicating that a sufficient spatial advantage has been gained. For lane-probing, the required constraints include longitudinal constraints for the following vehicle and ellipse constraints for the leading vehicle in the current group, in addition to lateral constraints. These ellipse constraints, described as linearized DHOCBF constraints in Section 3, enable the vehicle to safely probe forward for a spatial advantage while facilitating a smooth transition between lane-probing and lane-changing.
@@ V-A Constraints States
V-B Constraints Formulation
As mentioned in Section V-A, different states are associated with specific constraints. This subsection elaborates on the different constraints formulation, including longitudinal constraints, lateral constraints, ellipse constraints, safety boundary constraints, and actuator constraints.
Longitudinal Constraints. To ensure the safety of the ego vehicle longitudinally, it is crucial to maintain an appropriate distance from both the leading and following vehicles. Also, during lane-changing, longitudinal constraints on the desired group’s leader help ensure smooth adaptation. To achieve this, we adopt a linear DCBF aimed at defining a safety distance barrier. The function is mathematically formulated as
$h_{lon}(\mathbf{x})=|s-s_{i}|-T_{d}v_{x}-d_{0}$ (21)
where $s_{i}$ indicates the spacing of the surrounding vehicle with $i\in\{cl,cf,dl\}$. To be specific, $cl$ stands for the leader of the current group, $cf$ represents a follower in the current group, and $dl$ is the leader of the desired group. $T_{d}$ and $d_{0}$ denote the time headway and the least distance, respectively. The longitudinal constraints can hence be constructed by (15) as follows
$\Delta h_{lon}^{i}(\mathbf{x}_{k},\mathbf{u}_{k})\geq-\gamma_{lon}^{i}h_{lon}^{i}(\mathbf{x}_{k})-\epsilon_{lon}^{i}$ (22)
@@ V-B Constraints Formulation
$h_{lat}^{j}(\mathbf{x})=|e_{y}-e_{y}^{proj}|-w-d_{lat},\quad j\in\mathcal{J}$ (23)
where $\mathcal{J}=\{j\mid e_{y}^{\text{proj}}\in\text{ROI}\}$ $e_{y}^{proj}$ denotes the deviation distance projected on the track centerline of the ego vehicle. Moreover, $w$ is the width of a vehicle, and $d_{lat}$ is a positive constant. The lateral constraints can then be formed the same way as (15):
$\Delta h_{lat}^{j}(\mathbf{x}_{k},\mathbf{u}_{k})\geq-\gamma_{lat}^{j}h_{lat}^{j}(\mathbf{x}_{k})-\epsilon_{lat}^{j}$ (24)
@@
Ellipse Constraint. During lane-changing, we adopt linearized DHOCBF to formulate the constraint of the leader in the current vehicle group or the follower in the desired vehicle group. Different from [53], which used it for static obstacle avoidance and regarded obstacles as a circle, we apply it to dynamic traffic scenarios and adopt elliptical obstacles in the Frenet coordinate system.
@@
where the $s_{o}$ and $e_{yo}$ represent the traveled distance and lateral error of the considered ellipse center, respectively. $(s_{b},e_{yb})$ refers to the initial guess for the boundary of the ellipse.
@@
$L(s,e_{y},\lambda)=(s-s_{b})^{2}+(e_{y}-e_{y_{b}})^{2}$
$+\lambda\left(\frac{(s-s_{o})^{2}}{a^{2}}+\frac{(e_{y}-e_{y_{o}})^{2}}{b^{2}}-1\right)$ (26)
@@
The linearized DHOCBF is the tangent of the ellipse based on the closest point on the ellipse, the detailed formulation is as follows
@@
where the ellipse is centered at the front edge of the follower in the desired group and the rear edge of the leader in the current group. As illustrated in Fig. 6, the ego vehicle formulates individual ellipse constraints for both vehicles during the LC state and only for the follower during the LP state. These constraints are jointly incorporated into the optimization to ensure safe interactions.
Figure 6: Constraint mechanisms under different constraint states.
In the LK state, the optimization problem includes longitudinal DCBF constraints for the current lane and lateral DCBF constraints for adjacent lanes. In the LP state, the leading vehicle in the current lane is transformed into a linearized DHOCBF constraint to better accommodate potential lane change maneuvers while actively exploring the feasible space. In the LC state, the rear vehicle in the desired group and the front vehicle in the current group are the main interaction targets. Linearized DHOCBF ensures lane change safety and expands the feasible region for lane changes.
@@
$-\lambda_{b}-\epsilon_{b}\leq e_{y}\leq\lambda_{b}+\epsilon_{b}$ (28)
where $\lambda_{b}$ is a small positive constant and $\epsilon_{b}$ is the slack variable for safety boundary constraints. Safety boundary constraints restrict the lateral error in a certain range centered on the centerline. If the ego vehicle goes far from the centerline, its cost increases due to a penalty on $\epsilon_{b}$. In short, we describe this constraint as follows:
$H_{S}(\mathbf{x})\leq\epsilon_{b}$ (29)
@@
V-C Motion Planner Formulations
We employed Frenet-based MPC [55] to establish a finite-time convex optimization problem. Based on three different constraint states, we offer three optimization formulations. For simplicity, we abbreviate the three constraint states as LK (Lane-keeping), LP (Lane-probing), and LC (Lane-changing) optimizations. Fig. 6 provides a concise visual representation of the constraints associated with each of these states, illustrating how the optimization problem adapts to different driving conditions.
@@ V-C Motion Planner Formulations
$\underset{\mathbf{u}_{t}:t+N-1|t,\mathbf{x}_{t}:t+N|t}{\arg\min}\sum_{k=0}^{N-1}||\mathbf{u}_{t+k}||_{Q}+||\dot{\mathbf{u}}_{t+k}||_{P}$
@@
In the LK state, the optimization constraints include longitudinal constraints for the leader $\Delta h_{lon}^{cl}$ and follower $\Delta h_{lon}^{cf}$ vehicles at the current node, as well as lateral constraints $\Delta h_{lat}^{\mathcal{J}}$ for surrounding vehicles within the ROI.
@@
$\underset{\mathbf{u}_{t}:t+N-1|t,\mathbf{x}_{t}:t+N|t}{\arg\min}\sum_{k=0}^{N-1}||\mathbf{u}_{t+k}||_{Q}+||\dot{\mathbf{u}}_{t+k}||_{P}$
@@
In the LP state, the optimization constraints include longitudinal constraints for the follower $\Delta h_{lon}^{cf}$ at the current node, lateral constraints $\Delta h_{lat}^{\mathcal{J}}$ for surrounding vehicles within the ROI. Different from the LK state, the constraints for the leader vehicle at the current node are ellipse constraints, $\psi_{0}^{cl}$ and $\psi_{1}^{cl}$.
@@
$\underset{\mathbf{u}_{t}:t+N-1|t,\mathbf{x}_{t}:t+N|t}{\arg\min}\sum_{k=0}^{N-1}||\mathbf{u}_{t+k}||_{Q}+||\dot{\mathbf{u}}_{t+k}||_{P}$
@@
In the LC state, the optimization constraints include longitudinal constraints for the desired leader $\Delta h_{lon}^{dl}$ and current folower $\Delta h_{lon}^{cf}$, as well as ellipse constraints for both the desired follower $\psi_{0}^{df},\psi_{1}^{df}$ and the current leader $\psi_{0}^{cl},\psi_{1}^{cl}$, represented by the corresponding ellipse functions. Additionally, lateral constraints $\Delta h_{lat}^{\mathcal{J}}$ are applied to surrounding vehicles within the ROI to ensure safe lateral separation.
Among this three optimization formulation, (31b), (32b), (33b) denote the constraints imposed by system dynamics.
In the objectives in these three optimization formulation, the vector $\mathbf{x}_{ref}$ represents the desired values for $v_{x}$, $w$, $v_{y}$, along with the desired values for $e_{y}$ and $e_{\psi}$, which are set to 00. This setting arises because, in Frenet coordinates, all points are projected onto the centerline that the ego vehicle is intended to follow, thereby making the desired deviation distance and heading angle error equal to zero. $p(\mathbf{x}_{t+N_{p}|t})$ is the terminal cost, where $S$ signifies the weight attributed to the terminal cost. Drawing upon the method described by [22] that employs a control Lyapunov function (CLF) as the terminal cost, we utilize the squared error $e_{\psi}^{2}$ as the terminal cost. $||\mathbf{u}_{t+k}||$, $||\dot{\mathbf{u}}_{t+k}||$, and $||\mathbf{x}_{t+k}-\mathbf{x}_{ref}||$ serve as stage costs, with $Q$, $P$, and $R$ representing the weights for the control inputs, input derivatives, and quadratic terms of tracking $\mathbf{x}_{\text{ref}}$, respectively. In particular, the weight matrix $R$ for $v_{x}$, $e_{\psi}$, and $e_{y}$ is designed to increase over time. This time-increasing strategy helps to prevent the generation of excessively large steering angles at the beginning of the trajectory.
VI Experiment
In this section, we perform comprehensive simulations to evaluate the proposed system IDEAM. In section VI-A, we introduce the simulation background about scenarios and parameters. In section VI-B, we present the simulation benchmark including traditional methods and speed-oriented approaches, and evaluation metrics. Section VI-C gives the simulation results, and section VI-D shows a detailed case study of simulations.
VI-A Scenario and Parameters Description
@@
Simulation Scenarios. The simulation scenario occurs on a continuous closed-loop path that includes both curves and straight sections, challenging the autonomous system to maintain consistent performance across the varying road geometry. The surrounding vehicles are randomly generated within the scene, with varying initial positions and desired velocities, ensuring a diverse set of interactions. The longitudinal motion of surrounding vehicles is simulated using the intelligent driver model (IDM) [56], while lateral control is achieved through proportional-integral-derivative (PID) controller, the entire scenario takes place within a multi-lane environment, where the curvature of the bends varies along the path, reaching a maximum of approximately $0.1$. The dynamics of the scenario are updated at each step with a discretization interval of $\Delta t=100$ $\mathrm{ms}$, following a kinematic model that dictates vehicle movements as outlined below,
$\dot{x}$ $=v\cos(\psi+\beta)$ (34a)
$\dot{y}$ $=v\sin(\psi+\beta)$ (34b)
$\dot{\psi}$ $=\frac{v}{l_{r}}\sin\beta$ (34c)
$\dot{v}$ $=a$ (34d)
$\beta$ $=\tan^{-1}\left(\frac{l_{r}}{l_{f}+l_{r}}\tan\delta_{f}\right)$ (34e)